OpenStudy (anonymous):

determine whether convergent of divergent and evaluate integral of (x-1)^(-2/3)dx from 0 to 9

OpenStudy (anonymous):

\[\int\limits_{0}^{9}1/ \sqrt[3]{x^2-1}dx\] thT IS THE eq

OpenStudy (helder_edwin):

actually it is \[\large \int_0^9\frac{1}{(x-1)^{2/3}}\,dx \]

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

the function has a discontinuity at x=1. split the integral into two integral 0 to 1 and 1 to 9. and use limits. got it?

OpenStudy (anonymous):

ok got it does it converge

OpenStudy (helder_edwin):

i don't know. do the math.

OpenStudy (helder_edwin):

i think it does converge

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

u should do \[\large \int_0^9(x-1)^{-2/3}\,dx=\int_0^1(x-1)^{-2/3}\,dx+ \int_1^9(x-1)^{-2/3}\,dx \]

OpenStudy (helder_edwin):

\[\large \int_0^1(x-1)^{-2/3}\,dx=\lim_{b\to0+}\int_0^{1-b}(x-1)^{-2/3}\,dx\]

OpenStudy (helder_edwin):

do the same with the other with 1+b instead of 1-b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

why b-1

OpenStudy (helder_edwin):

because u have to approach 1 from the left, since u r on the interval [0,1]. it is 1-b, NOT b-1